Olivia, a noncorporate taxpayer, experienced the following Sec. 1231 gains and losses during the years 2015 through 2018. Her first disposition of a Sec. 1231 asset occurred in 2015. What is the tax treatment for 2018?

gains losses
2015 18000 23000
2016 9000 13000
2017 22000 16000
2018 25000 17000

A) $8,000 LTCG

B) $8,000 Ordinary Income

C) $4,000 Ordinary Income, $4,000 LTCG

D) $3,000 Ordinary Income, $5,000 LTCG
Answer #1

D. $3000 ordinary income, $5000 LTCG

As per IRS Provision applied : Nonrecaptured section 1231 losses are net section 1231 losses for the previous 5 years that have not been applied against a net section 1231 gain. So, if in any of your 5 preceding tax years you had section 1231 losses, a net gain for the current year from the sale of section 1231 assets is ordinary gain to the extent of your prior losses. The amount above the prior losses shall be treated as long term capital gain

The net losses is calculated below:

2015 = 18000 -23000 = $5000 Loss

2016 = 9000 – 13000 = $4000 Loss

2017 = 22000 – 16000 = $6000 Gain

2018 = 25000 – 17000 = $8000 Gain

Net section 1231 gain

8000

Net section 1231 loss 2015

(5000)

Net section 1231 loss 2016

(4000)

Net section 1231 gain 2017

6000

Remaining net section 1231 losses from prior 5 years

(3000)

Gain treated as ordinary income

$3000

Gain treated as long term capital gain

$5000

Kindly upvote